# UniCUE: Unified Recognition and Generation Framework for Chinese Cued Speech Video-to-Speech Generation

### v1

@@ 4.6. User Study
To holistically evaluate the perceptual quality of synthesized speech, we conduct a user study. In this experiment, test samples were randomly selected from the outputs of each cuer, and 10 volunteers were invited to rate the generated speech in three perceptual dimensions on 5-point Likert scales: Accuracy (1: Unintelligible, 5: Perfect), Quality (1: Artificial, 5: Human-like), and Synchronization (1: Desynchronized, 5: Perfectly aligned). Four systems were compared: the SOTA CSR baseline (CMML (Liu and Liu, 2023)), our CSR, CSV2S, and the proposed UniCUE. Based on the mean and variance values presented in Figure 8, UniCUE demonstrates statistically significant superiority across all metrics. These results confirm that the unified framework successfully bridges the gap between visual understanding and speech generation, outperforming both combined pipelines and task-specific baselines in human perception.
4.7. Discussion and Analysis
Analysis on Cuer Representation. To quantitatively assess the capability in preserving cuer-specific acoustic characteristics, we analyze the latent cuer embeddings extracted from generated speech using Resemblyzer666https://github.com/resemble-ai/Resemblyzer. The embeddings are projected into 2D space via t-SNE (Maaten and Hinton, 2008) for visual interpretation. As shown in Figure 9, speech samples from the same cuer (e.g., cuer 1 to cuer 6) form tight cluster in the high-dimensional embedding space, while the distances between the cluster centers of different cuers are significantly larger. This indicates that the model effectively captures and distinguishes the personalized acoustic features of each cuer.
Computational Efficiency. To quantify the efficiency gains of our unified framework, we conduct time cost evaluations across both training and inference stage. (a) Training Efficiency. We compared the training performance of CSV2S and UniCUE to evaluate the training efficiency of our proposed unified framework. As shown in the Figure 10, the training loss of UniCUE decreases rapidly to a stable value in fewer epochs, indicating that the introduction of fine-grained understanding ability accelerates the learning process of the speech generation model. Moreover, the validation loss also decreases concurrently with less fluctuation, demonstrating improved model robustness. (b) Real-time Inference. To validate the real-time performance of the model, this study further compares the inference efficiency of combined(our CSR and TTS) and unified CSV2S. As an end-to-end CSV2S generation framework, our UniCUE significantly reduces the inference time (4.15s), achieving a 40% reduction compared to the traditional CSR + TTS serial approach (6.93s), thereby confirming the advantages of the unified architecture in efficiency optimization.
@@ 5. Conclusion
This work introduces UniCUE, the first unified framework that integrates Cued Speech recognition (CSR) and video-to-speech (CSV2S) generation for Chinese Mandarin CS, addressing the critical challenges of bridging communication barriers between the hearing-impaired and normal-hearing communities. By integrating CS recognition with diffusion-based speech generation, our framework achieves fine-grained visual-semantic alignment and precise temporal synchronization, even with limited CS training data. Key innovations, such as the pose-aware visual processor for detailed visual embedding, the semantic alignment pool for robust visual-text mapping, and the VPA for injecting CSR knowledge into the diffusion decoder, collectively enable speech synthesis that preserves both linguistic accuracy and temporal consistency. Experimental results demonstrate that our unified approach not only outperforms SOTA CSR methods in semantic consistency but also achieves strong performance in temporal alignment. Furthermore, the expanded Chinese CS dataset establishes a solid foundation for future research.
@@ Acknowledgements.
- Chen et al. (2025a) Liang Chen, Shuai Bai, Wenhao Chai, Weichu Xie, Haozhe Zhao, Leon Vinci, Junyang Lin, and Baobao Chang. 2025a. Multimodal Representation Alignment for Image Generation: Text-Image Interleaved Control Is Easier Than You Think. arXiv:2502.20172 (2025).
- Chen et al. (2025b) Xiaokang Chen, Zhiyu Wu, Xingchao Liu, Zizheng Pan, Wen Liu, Zhenda Xie, Xingkai Yu, and Chong Ruan. 2025b. Janus-Pro: Unified Multimodal Understanding and Generation with Data and Model Scaling. arXiv preprint arXiv:2501.17811 (2025).
- Chen et al. (2024) Yushen Chen, Zhikang Niu, Ziyang Ma, Keqi Deng, Chunhui Wang, Jian Zhao, Kai Yu, and Xie Chen. 2024. F5-tts: A fairytaler that fakes fluent and faithful speech with flow matching. arXiv:2410.06885 (2024).
- Choromanski et al. (2020) Krzysztof Choromanski, Valerii Likhosherstov, David Dohan, Xingyou Song, Andreea Gane, Tamas Sarlos, Peter Hawkins, Jared Davis, Afroz Mohiuddin, Lukasz Kaiser, et al. 2020. Rethinking attention with performers. arXiv:2009.14794 (2020).
@@
- Dong et al. (2023) Runpei Dong, Chunrui Han, Yuang Peng, Zekun Qi, Zheng Ge, Jinrong Yang, Liang Zhao, Jianjian Sun, Hongyu Zhou, Haoran Wei, et al. 2023. Dreamllm: Synergistic multimodal comprehension and creation. arXiv:2309.11499 (2023).
- Fang et al. (2024) Rongyao Fang, Chengqi Duan, Kun Wang, Hao Li, Hao Tian, Xingyu Zeng, Rui Zhao, Jifeng Dai, Hongsheng Li, and Xihui Liu. 2024. Puma: Empowering unified mllm with multi-granular visual generation. arXiv:2410.13861 (2024).
- Ge et al. (2024a) Yuying Ge, Yizhuo Li, Yixiao Ge, and Ying Shan. 2024a. Divot: Diffusion Powers Video Tokenizer for Comprehension and Generation. arXiv:2412.04432 (2024).
- Ge et al. (2024b) Yuying Ge, Sijie Zhao, Jinguo Zhu, Yixiao Ge, Kun Yi, Lin Song, Chen Li, Xiaohan Ding, and Ying Shan. 2024b. Seed-x: Multimodal models with unified multi-granularity comprehension and generation. arXiv:2404.14396 (2024).
@@
- Hershey et al. (2017) Shawn Hershey, Sourish Chaudhuri, Daniel PW Ellis, Jort F Gemmeke, Aren Jansen, R Channing Moore, Manoj Plakal, Devin Platt, Rif A Saurous, Bryan Seybold, et al. 2017. CNN architectures for large-scale audio classification. In 2017 IEEE international conference on acoustics, speech and signal processing (ICASSP). IEEE, 131–135.
- Hua et al. (2022) Weizhe Hua, Zihang Dai, Hanxiao Liu, and Quoc Le. 2022. Transformer quality in linear time. In International conference on machine learning. PMLR, 9099–9117.
- Jin et al. (2023) Yang Jin, Kun Xu, Liwei Chen, Chao Liao, Jianchao Tan, Bin Chen, Chenyi Lei, An Liu, Chengru Song, Xiaoqiang Lei, et al. 2023. Unified language-vision pretraining with dynamic discrete visual tokenization. arXiv:2309.04669 (2023).
@@
- Kilgour et al. (2018) Kevin Kilgour, Mauricio Zuluaga, Dominik Roblek, and Matthew Sharifi. 2018. Fr’echet audio distance: A metric for evaluating music enhancement algorithms. arXiv:1812.08466 (2018).
@@
- Lei et al. (2024b) Wentao Lei, Li Liu, and Jun Wang. 2024b. Bridge to Non-Barrier Communication: Gloss-Prompted Fine-grained Cued Speech Gesture Generation with Diffusion Model. arXiv:2404.19277 (2024).
@@
- Li et al. (2024) Yanwei Li, Yuechen Zhang, Chengyao Wang, Zhisheng Zhong, Yixin Chen, Ruihang Chu, Shaoteng Liu, and Jiaya Jia. 2024. Mini-gemini: Mining the potential of multi-modality vision language models. arXiv:2403.18814 (2024).
- Liu et al. (2023) Haohe Liu, Zehua Chen, Yi Yuan, Xinhao Mei, Xubo Liu, Danilo Mandic, Wenwu Wang, and Mark D Plumbley. 2023. Audioldm: Text-to-audio generation with latent diffusion models. arXiv:2301.12503 (2023).
@@
- Liu et al. (2020) Li Liu, Gang Feng, Denis Beautemps, and Xiao-Ping Zhang. 2020. Re-synchronization using the hand preceding model for multi-modal fusion in automatic continuous cued speech recognition. IEEE Transactions on Multimedia 23 (2020), 292–305.
@@
- Liu (2020) Y Liu. 2020. Multilingual denoising pre-training for neural machine translation. arXiv:2001.08210 (2020).
- Maaten and Hinton (2008) Laurens van der Maaten and Geoffrey Hinton. 2008. Visualizing data using t-SNE. Journal of Machine Learning Research (JMLR) 9, 86 (2008), 2579–2605. http://www.jmlr.org/papers/v9/vandermaaten08a.html
@@
- Rong and Liu (2024) Yan Rong and Li Liu. 2024. Seeing your speech style: A novel zero-shot identity-disentanglement face-based voice conversion. arXiv preprint arXiv:2409.00700 (2024).
@@
- Wang et al. (2020) Sinong Wang, Belinda Z Li, Madian Khabsa, Han Fang, and Hao Ma. 2020. Linformer: Self-attention with linear complexity. arXiv:2006.04768 (2020).
- Wu et al. (2024a) Chengyue Wu, Xiaokang Chen, Zhiyu Wu, Yiyang Ma, Xingchao Liu, Zizheng Pan, Wen Liu, Zhenda Xie, Xingkai Yu, Chong Ruan, et al. 2024a. Janus: Decoupling visual encoding for unified multimodal understanding and generation. arXiv:2410.13848 (2024).
- Wu et al. (2023) Shengqiong Wu, Hao Fei, Leigang Qu, Wei Ji, and Tat-Seng Chua. 2023. Next-gpt: Any-to-any multimodal llm. arXiv:2309.05519 (2023).
@@
- Yemini et al. (2023) Yochai Yemini, Aviv Shamsian, Lior Bracha, Sharon Gannot, and Ethan Fetaya. 2023. LipVoicer: Generating Speech from Silent Videos Guided by Lip Reading. arXiv:2306.03258 (2023).
@@
- Zhen et al. (2022) Q Zhen, W Sun, H Deng, D Li, Y Wei, B Lv, J Yan, L Kong, and Y Zhong. 2022. cosFormer: Rethinking softmax in attention. In International Conference on Learning Representations.
- Zhou et al. (2024) Chunting Zhou, Lili Yu, Arun Babu, Kushal Tirumala, Michihiro Yasunaga, Leonid Shamis, Jacob Kahn, Xuezhe Ma, Luke Zettlemoyer, and Omer Levy. 2024. Transfusion: Predict the next token and diffuse images with one multi-modal model. arXiv:2408.11039 (2024).

### v3

@@ 4.4. User Study
To comprehensively assess the perceptual quality of synthesized speech, we conduct a user study involving 20 randomly selected test samples per cuer. Twenty volunteers rate the generated speech on three perceptual dimensions using 5-point Likert scales: Accuracy (1: unintelligible, 5: perfectly intelligible), Quality (1: artificial, 5: human-like), and Synchronization (1: desynchronized, 5: perfectly aligned). As shown in Figure 5, UniCUE consistently achieves significantly higher scores across all metrics, demonstrating statistically meaningful improvements. These findings validate that our unified framework effectively bridges visual understanding and speech generation, delivering superior performance in human perception compared to both modular pipelines and task-specific baselines.
@@ 5. Conclusion
This work introduces UniCUE, the first unified framework for directly generating speech from CS videos. By integrating fine-grained visual understanding with diffusion-based speech synthesis, UniCUE produces intelligible speech with precise temporal alignment. Key components including the pose-aware visual processor, semantic alignment pool, and VisioPhonetic Adapter, enable effective knowledge transfer from CS recognition (CSR) to CS video-to-speech generation (CSV2S), enhancing both linguistic accuracy and temporal synchronization. Additionally, we introduce the UniCUE-HI corpus, a new CS dataset featuring both normal-hearing and hearing-impaired cuers. Extensive experiments on this dataset demonstrate that UniCUE state-of-the-art methods across multiple evaluation metrics.
@@ Acknowledgements.
- Chen et al. (2025a) Liang Chen, Shuai Bai, Wenhao Chai, Weichu Xie, Haozhe Zhao, Leon Vinci, Junyang Lin, and Baobao Chang. 2025a. Multimodal Representation Alignment for Image Generation: Text-Image Interleaved Control Is Easier Than You Think. CoRR (2025).
- Chen et al. (2025b) Xiaokang Chen, Zhiyu Wu, Xingchao Liu, Zizheng Pan, Wen Liu, Zhenda Xie, Xingkai Yu, and Chong Ruan. 2025b. Janus-Pro: Unified Multimodal Understanding and Generation with Data and Model Scaling. CoRR (2025).
- Choi et al. (2023) Jeongsoo Choi, Minsu Kim, and Yong Man Ro. 2023. Intelligible Lip-to-Speech Synthesis with Speech Units. In Proceedings of the Annual Conference of the International Speech Communication Association, INTERSPEECH, Vol. 2023. 4349–4353.
@@
- Dong et al. (2024) Runpei Dong, Chunrui Han, Yuang Peng, Zekun Qi, Zheng Ge, Jinrong Yang, Liang Zhao, Jianjian Sun, Hongyu Zhou, Haoran Wei, et al. 2024. DreamLLM: Synergistic Multimodal Comprehension and Creation. In ICLR.
- Fang et al. (2024) Rongyao Fang, Chengqi Duan, Kun Wang, Hao Li, Hao Tian, Xingyu Zeng, Rui Zhao, Jifeng Dai, Hongsheng Li, and Xihui Liu. 2024. PUMA: Empowering Unified MLLM with Multi-granular Visual Generation. CoRR (2024).
- Ge et al. (2025) Yuying Ge, Yizhuo Li, Yixiao Ge, and Ying Shan. 2025. Divot: Diffusion powers video tokenizer for comprehension and generation. In Proceedings of the Computer Vision and Pattern Recognition Conference. 13606–13617.
- Ge et al. (2024) Yuying Ge, Sijie Zhao, Jinguo Zhu, Yixiao Ge, Kun Yi, Lin Song, Chen Li, Xiaohan Ding, and Ying Shan. 2024. SEED-X: Multimodal Models with Unified Multi-granularity Comprehension and Generation. CoRR (2024).
- Graves et al. (2006) Alex Graves, Santiago Fernández, Faustino Gomez, and Jürgen Schmidhuber. 2006. Connectionist temporal classification: labelling unsegmented sequence data with recurrent neural networks. In Proceedings of the 23rd international conference on Machine learning. ACM, 369–376.
@@
- Jin et al. (2024) Yang Jin, Kun Xu, Liwei Chen, Chao Liao, Jianchao Tan, Quzhe Huang, Bin Chen, Chengru Song, Dai Meng, Di Zhang, et al. 2024. Unified Language-Vision Pretraining in LLM with Dynamic Discrete Visual Tokenization. In ICLR.
@@
- Lei et al. (2024b) Wentao Lei, Li Liu, and Jun Wang. 2024b. Bridge to non-barrier communication: gloss-prompted fine-grained cued speech gesture generation with diffusion model. In Proceedings of the Thirty-Third International Joint Conference on Artificial Intelligence. 6333–6341.
@@
- Li et al. (2024) Yanwei Li, Yuechen Zhang, Chengyao Wang, Zhisheng Zhong, Yixin Chen, Ruihang Chu, Shaoteng Liu, and Jiaya Jia. 2024. Mini-Gemini: Mining the Potential of Multi-modality Vision Language Models. CoRR (2024).
- Liu et al. (2023) Haohe Liu, Zehua Chen, Yi Yuan, Xinhao Mei, Xubo Liu, Danilo Mandic, Wenwu Wang, and Mark D Plumbley. 2023. AudioLDM: Text-to-Audio Generation with Latent Diffusion Models. In International Conference on Machine Learning. PMLR, 21450–21474.
@@
- Liu et al. (2020a) Li Liu, Gang Feng, Denis Beautemps, and Xiao-Ping Zhang. 2020a. Re-synchronization using the hand preceding model for multi-modal fusion in automatic continuous cued speech recognition. IEEE Transactions on Multimedia 23 (2020), 292–305.
@@
- Liu et al. (2020b) Yinhan Liu, Jiatao Gu, Naman Goyal, Xian Li, Sergey Edunov, Marjan Ghazvininejad, Mike Lewis, and Luke Zettlemoyer. 2020b. Multilingual denoising pre-training for neural machine translation. Transactions of the Association for Computational Linguistics 8 (2020), 726–742.
@@
- Reddy et al. (2021) Chandan KA Reddy, Vishak Gopal, and Ross Cutler. 2021. DNSMOS: A non-intrusive perceptual objective speech quality metric to evaluate noise suppressors. In ICASSP 2021-2021 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP). IEEE, 6493–6497.
@@
- Taal et al. (2010) Cees H Taal, Richard C Hendriks, Richard Heusdens, and Jesper Jensen. 2010. A short-time objective intelligibility measure for time-frequency weighted noisy speech. In 2010 IEEE international conference on acoustics, speech and signal processing. IEEE, 4214–4217.
- Wu et al. (2025) Chengyue Wu, Xiaokang Chen, Zhiyu Wu, Yiyang Ma, Xingchao Liu, Zizheng Pan, Wen Liu, Zhenda Xie, Xingkai Yu, Chong Ruan, et al. 2025. Janus: Decoupling visual encoding for unified multimodal understanding and generation. In Proceedings of the Computer Vision and Pattern Recognition Conference. 12966–12977.
- Wu et al. (2024a) Shengqiong Wu, Hao Fei, Leigang Qu, Wei Ji, and Tat-Seng Chua. 2024a. NExT-GPT: Any-to-Any Multimodal LLM. In International Conference on Machine Learning. PMLR, 53366–53397.
@@
- Yemini et al. (2024) Yochai Yemini, Aviv Shamsian, Lior Bracha, Sharon Gannot, and Ethan Fetaya. 2024. LipVoicer: Generating Speech from Silent Videos Guided by Lip Reading. In The Twelfth International Conference on Learning Representations.
@@
- Zhou et al. (2024) Chunting Zhou, LILI YU, Arun Babu, Kushal Tirumala, Michihiro Yasunaga, Leonid Shamis, Jacob Kahn, Xuezhe Ma, Luke Zettlemoyer, and Omer Levy. 2024. Transfusion: Predict the Next Token and Diffuse Images with One Multi-Modal Model. In The Thirteenth International Conference on Learning Representations.